### v1

# Dracula’s Chivito: discovery of a large edge-on protoplanetary disk with Pan-STARRS

Abstract
We report the serendipitous discovery of a large edge-on protoplanetary disk associated with the infrared source IRAS 23077+6707. The disk’s apparent size in the Pan-STARRS (PS1) images is , making this one of the largest known disks on the sky. It is likely a young system, still surrounded by the envelope which is very faint but still visible in the PS1 images in the northern part (alternatively this structure could be filaments from the disk itself). We use the PS1 magnitudes and other available photometric data to construct the spectral energy distribution (SED) of the disk. An optical spectrum indicates that the obscured star is hot, most likely late A. We adopt a distance of 300 pc for this object based on Gaia DR3 extinctions. We model the system using the HOCHUNK3D radiative transfer software and find that the system is consistent with a hot star of effective temperature 8000 K surrounded by a disk of size 1650 AU and mass 0.2 M at inclination 82.
1 Introduction
Planet formation within protoplanetary disks around ostensibly similar stars results in a surprisingly large diversity of exoplanetary systems (e.g., Winn & Fabrycky, 2015, and references therein). The processes responsible involve the hierarchical assemblage of successively larger bodies, beginning with sub-micron-sized dust particles, but the details of these processes and how they lead to the diverse exoplanet demographics revealed by the intensive planet hunting endeavours over the last two decades remain poorly understood. Key to this is dust growth and transport, and the extent to which the dust couples to the gas as a function of grain size, gas temperature and turbulence (Goldreich & Ward, 1973; Weidenschilling, 1977; Voelk et al., 1980; Dullemond & Dominik, 2005; Binkert, 2023).
Extensive studies of disks aimed at making progress in this direction have tended to focus on low-inclination objects in which remarkable structures such as spirals, rings and gaps have been resolved (e.g., ALMA Partnership et al., 2015; Ruge et al., 2016; Andrews, 2020). However, it is difficult to determine the vertical structure of these disks (e.g., Pinte et al., 2016), the relative distributions of gas and dust, and the radial dependence and extent of dust settling and disk flaring. It has been increasingly recognised in recent years that edge-on protoplanetary disks instead offer the opportunity to probe the three-dimensional morphology of disks, sacrificing disk plane details but providing a direct line of sight of the vertical axis. The shadows cast by the midplane concentration of dust and gas in highly-inclined disks allow direct observations of scattered light from the central star, unveiling the vertical distribution of dust particles, providing information about their sizes, compositions, and settling mechanisms (Dutrey et al., 2017; Villenave et al., 2020; Wolff et al., 2021; Angelo et al., 2023). This different perspective into disk structure and evolution renders nearby edge-on disks in which spatial resolution is maximized of particularly great value. In this Letter, we report the serendipitous discovery of a remarkable new edge-on protoplanetary disk candidate using images taken by the Panoramic Survey Telescope and Rapid Response System (Pan-STARRS, hereafter PS1). We shall argue that the central star is an obscured object of intermediate mass and likely a Herbig Ae star, and that distance constraints imply that the disk is possibly the largest protoplanetary disk found to date.
The structure of the paper is as follows. Section 2 describes the discovery. In Section 3 we present the photometric data used to construct the SED and the spectroscopic data to constrain the host star properties. In Section 4 we estimate the distance to this object. In Section 5 we model the disk using the HOCHUNK3D radiative transfer code. In Section 6 we discuss our results and in Section 7 we summarize our conclusions.
2 Discovery of DraChi
This remarkable disk was discovered in PS1 images while working on a variability study of active galactic nuclei candidates from Secrest et al. (2015) in conjunction with data from the USNO Robotic Astrometric Telescope (Zacharias et al., 2015). PS1 observed the entire sky north of declination in five filters (g, r, i, z, y, hereafter ). A composite color PS1 image of the disk is illustrated in the left panel of Figure 1. The disk has the characteristic bipolar appearance of edge-on disks with the central star being completely obscured (Figure 1). The bright emission covers about on the sky, with a very faint structure in the disk’s northern part extending out to 18. DraChi appears blue in comparison with the stars in the vicinity suggesting a hot star behind the disk.
The object was not unknown: it was observed earlier at radio wavelengths as a possible pre-main sequence star (PMS) by Wouterloot et al. (1993). More recently, an optical spectrum was obtained by a French amateur astronomy group looking for planetary nebula (PNe) candidates111https://planetarynebulae.net. They named the disk Kn 32 in their PNe sample..
There are very few edge-on disks known so far, especially such large disks, making this one particularly valuable for study and future observations. It is not associated with any known star forming region, similar to the well-known Gomez’s Hamburger (“GoHam”, Ruiz et al., 1987), in size. Submillimeter Array observations revealed Keplerian rotation about the central star and established GoHam as a massive circumstellar disk (Bujarrabal et al., 2008; Wood et al., 2008). Subsequent observations constrained the distance and the disk and stellar mass estimates to pc, , and , respectively (Bujarrabal et al., 2008; Berné et al., 2015), and indicated the possible presence of a protoplanet (Berné et al., 2015; Teague et al., 2020).
Since our disk is similar to GoHam, and in addition has faint “fangs” extending far out from the northern edges (see Figure 1), we name it “Dracula’s Chivito”222One of the authors grew up in Transylvania, not far from the place where Vlad (Draculea) Tepes lived, and another author is from Uruguay, where “chivito” is a traditional sandwich and it is similar to a hamburger., DraChi hereafter. Other similar objects are PDS 144N (Perrin et al., 2006; Hornbeck et al., 2012), PDS 453 (Perrin et al., 2010), IRAS 04302+2247 (Padgett et al., 1999), ESO-H 569 (Wolff et al., 2017), HH 30 (Burrows et al., 1996), HK Tau/c (Stapelfeldt et al., 1998), HV Tau C (Monin & Bouvier, 2000) and SSTC2DJ163131.2–242627 (Wolff et al., 2021). However, DraChi is by far the largest of these, and at in angular extent is the largest edge-on protoplanetary disk discovered to date.
3 Photometry and Spectroscopy
Stacked PS1 images were retrieved for the filters centered on the sky location , . To perform the photometric analysis, we followed the procedure described in the PS1 archive333https://outerspace.stsci.edu/display/PANSTARRS/PS1+Stack+images and by Waters et al. (2020), using an aperture of 14 to cover the whole disk. There is a small amount of contamination from at least one faint star on the right side of the disk but it is insignificant for the purpose of our analysis.
We also searched for all the available archival photometric data on DraChi. These and our own photometry are shown in Table 1. As noted for PS1, some contamination from stars is expected in all these measurements, but it is likely very small, especially at longer wavelengths, where the disk dominates the nearby sources. We estimate that the contamination level is always less than 25%. We also note that some variability at this level is also expected for PMS stars (e.g. Espaillat et al., 2011). We use these data in Section 5 to construct the SED and constrain the simulated models.
The optical spectrum obtained by the French amateur astronomers was taken on Aug 3, 2019. More details about the exemplary work of this group can be found in Le Dû et al. (2022). Our object was named Kn 32 and is included in Table 4 of that work. Some details of the observation can be found on their website444https://planetarynebulae.net/EN/tableau_spectres.php. We fitted the spectrum to star templates based on spectra taken by the Sloan Digital Sky Survey using the software PyHammer v2.0 (https://github.com/BU-hammerTeam/PyHammer, Kesseli et al., 2017; Roulston et al., 2020). We find a best fit for a type A9 star and extinction (see Figure 2).
4 Distance
Since DraChi is not associated with any known star-forming region the distance cannot be precisely determined, similar to GoHam which has estimates ranging from 250 to 500 pc. However, new stars form in interstellar clouds so we can look for the cloud associated with DraChi based on extinctions. We used Gaia DR3 extinctions estimates for stars located within 3 distance from DraChi. The Gaia extinctions were converted to Av using the dustapprox tool555https://github.com/mfouesneau/dustapprox. The result is shown in Figure 3. The nearest cloud seems to be located between 70 and 130 pc with A up to 0.7 and the next larger cloud between 250 and 500 pc with A up to 1.4. Based on constraints obtained form the modeling in the next section we find that DraChi is unlikely to be located in the first cloud because the inferred star luminosity is too low for an A star. It is also unlikely to be located beyond 350 pc because of the apparent size of the disk and large absorption values past this distance. We adopt a distance of 300 pc for our modeling and in the discussion section we present more arguments for choosing this distance.
5 Radiation Transfer Modeling
To investigate very approximate disk parameters based on the various data at hand we use the three-dimensional dust continuum Monte-Carlo radiative transfer code HOCHUNK3D666https://gemelli.colorado.edu/ bwhitney/codes/codes.html (Whitney et al., 2013). The basic structure of a typical flared disk can be described by the height of the disk, the surface density, and the density structure, respectively:
where is the disk scale height exponent, is the disk radial density exponent, and and are the scale height and surface density of the disk at the pivot radius, .
We first use the optical spectrum and the PS1 images to impose some constraints on the parameters. The optical spectrum is consistent with a star temperature in the range 6500 - 8500 K. The PS1 images constrain the inclination angle to 80 - 84 and the apparent size of the disk to 11.
Finally, based on the distance of 300 pc estimated in Sect. 4, we can estimate the width of the disk and the scale height constrained by the PS1 images. We find a scale height of 0.05 - 0.1 at a radius of 500 AU. The dust destruction radius as defined in Robitaille et al. (2006) is given by , with the dust sublimation temperature . For approximate values of and , or 0.25 AU, which we set as the inner radius of the disk.
Miotello et al. (2014) describe Class I objects as disks with partially dissipating envelopes that contain large grains embedded deep within the inner regions of the envelopes. Such an envelope is suggested by the faint “fangs” seen in the PS1 images. Our model for DraChi includes a disk with large grains settled closer to the mid-plane as well as the usual small grain disk plus an Ulrich envelope (Ulrich, 1976).
HOCHUNK3D allows for two coplanar disks with their own parameters for the structure and dust composition. One disk contains large dust grains with a small scale height, and the other small dust grains with a larger scale height. Both disks begin at the sublimation radius of 0.25 AU. The small grain disk ends at 1650 AU, whereas the large grain disk ends at 500 AU which best fits the near-IR. For the large grains, we used Model 1 of Wood et al. (2001), which combines a power-law with exponential cutoff for components of amorphous carbon and astronomical silicates with solar abundance constraints for both, and a max radius of a = 1000 m. For the small grains, we used the model of Kim et al. (1994), which is a mixture of silicates and graphite where the size distribution is determined by maximum-entropy fitting of the interstellar extinction curve. For the silicates, a = 0.16 m and for the graphite, a = 0.28 m. Both disks also include polycyclic aromatic hydrocarbons and the very small grains (PAHs/VSGs) dust model from Draine & Li (2007). The Ulrich envelope has the same composition as the small grains disk, has an ambient density at the inner sublimation radius of g/s and extends to 3000 AU.
Exhaustive radiation transfer modelling is beyond the scope of this letter; the intention here is instead to estimate a plausible set of coarse disk parameters that explain the data reasonably well. Given the constraints imposed by the PS1 image and the A9 spectral type, we explored a range of parameters to find a model that best fits the SED and especially the far-IR as it is not affected by extinction. The small and large grains scale heights are 0.06 and 0.01 at 500 AU, respectively. The scale height exponent is 1.15 for the small grain disk and 1.4 for the large grains. The radial density for both disks was adjusted to best fit both near-IR and far-IR data, with = 2.2 for both disks. The dust-to-gas ratio was taken at the standard fraction of 0.01.
The best model is presented in Figures 1 and 4 which show the model image convolved with the PS1 PSFs, and the model SED, respectively. We emphasise that this model is not unique, but it provides a good match to the available data. However, there are significant discrepancies in the near and mid-IR range which we discuss below. It is also remarkable how well the envelope included in the model resembles the faint “fangs” seen in the PS1 image in the northern part of the disk (Figure 1). The stellar effective temperature is 8000 K with a luminosity 11.46 L. The total disk mass is 0.2 M, with 20% of it in the large grains. The inclination angle is 82. A line-of-sight extinction was added to the model.
6 Discussion
6.1 DraChi: the largest known protoplanetary disk
DraChi is a newly-identified edge-on protoplanetary disk and in angular extent is the largest such object now known. This identification is based on PS1 images, and the results of radiation transfer modeling are also consistent with a late A-type star surrounded by a protoplanetary disk seen almost edge-on, and by an almost dissipated envelope.
DraChi is very similar to the well known edge-on protoplanetary disk GoHam, they both likely have HAeBe central stars, have a large apparent size and are not associated with known star clusters. One significant difference is the presence of the faint “fangs” in the PS1 images of DraChi, which we interpret as a dissipating envelope. DraChi is about 50% larger than GoHam in angular extent. We note that GoHam presents evidence for a possible protoplanet (Berné et al., 2015), showing that planet formation might be detected in such edge-on disks.
6.2 The SED
Modelling of the SED shown in Figure 4 presents a good overall match to observations, but discrepancies occur in the near- and mid-IR. We included PAH/VSG emisison in an attempt to reproduce the SED shape following Wood et al. (2008), but there are still significant discrepancies in mid-infrared around 10-20 m. We note that the photometric data is not simultaneous and such disks are known to be variable. The central star luminosity and temperature match a PMS star of type late A.
6.3 The “Fangs”
The radiative transfer model also reproduces well the PS1 image (Figure 1), including the faint “fangs” extending 18 in the northern part of the disk. The “fangs” are very similar to the “edge” of the envelope created by the shadow of the disk on the illuminated envelope. The southern part of this envelope is likely obscured in the PS1 images but it might be detected with more sensitive observations.
Another possibility is that the “fangs” are filaments associated with the disk itself. Smaller such filaments can be seen on the left side of the disk but these are located close to the central part of the disk, suggesting a possible outflow. If indeed the envelope is still present, this suggests a young disk at the end of the Class I phase. Another significant difference is the pronounced flaring seen in both lobes but most clearly in the southern part of the right lobe. The flaring also makes the left lobe curve outward, in contrast to the model image. Increasing the flaring in our model could not reproduce these features. Such flaring was seen in a similar HAeBe edge-on disk, PDS 144N (Perrin et al., 2006), which was also modelled with an envelope.
6.4 The Distance to DraChi
Estimating the distance to edge-on disks is hard if they are not associated with known star-forming regions, as demonstrated by GoHam with estimates between 250 and 500 pc, or PDS 144N (Perrin et al., 2006). The same is true for DraChi, and here we present arguments for preferring the adopted distance of 300 pc as opposed to a distance smaller than 150 pc, corresponding to the closest cloud found in Section 4, or larger distances of 350 - 500 pc.
Firstly, for a distance of 300 pc the disk radius is 1650 AU, which is already at the higher end of the range of radii for known disks. For larger distances of 350 - 500 pc the disk radius should be 1925 - 2750 pc which is unlikely. Secondly, the extinction estimated from Gaia for distances corresponding to the nearest cloud in Figure 3 are smaller than 0.7 A, and those for the next cloud go up to 1.4 A. Both our estimates based on the star spectrum and the radiation transfer modeling are close to 0.9 A which is the extinction corresponding to 300 pc in Figure 3, indicating that DraChi is likely in the second cloud but at the front part of it. Finally, if DraChi is located in the nearest cloud the luminosity of the star should be too small for an A star. For 300 pc the modeling suggests a luminosity of 11.5 L, but for distances less than 150 pc the star luminosity should be less than 5.0 L which is unlikely for a HAeBe star (Montesinos et al., 2009). We should note, however, that our estimated stellar type is based on scattered light alone as the star is not seen directly and that the observed luminosity in edge-on disks depends on multiple factors such as flaring parameter, disk mass, etc. Finally, a similar argument is based on the mass of the disk estimated from the radiation transfer modeling. The disk mass is given mostly by the far-infrared peak seen on the SED so it is not affected much by extinction but is also determined by the disk geometry such as flaring. Our model implies a disk mass of . If the distance is larger, the disk mass estimate increases to over for a distance of 500 pc, which again becomes difficult to explain in the context of known protoplanetary disk parameters.
For all these reasons we prefer the distance of 300 pc, and consider a distance smaller than 150 pc or larger than 350 pc unlikely.
7 Conclusion
We have identified a new, very large protoplanetary disk from PS1 images and confirmed the consistency of available data with a disk interpretation through radiative transfer and SED modelling. The host star was estimated to be an A9 HAeBe star based on an archival optical spectrum. We use the PS1 image to constrain the radiation transfer model parameters and the SED to find the best model. While the data is insufficient to perform detailed modeling, our model matches the observed SED well, although discrepancies are present in the mid-IR. The model disk has a mass of 0.2 , an inclination angle of 82, and its radius is 1650 AU at the adopted distance of 300 pc estimated from Gaia stellar extinctions. The low-density envelope included in the model has a 3000 AU radius and reproduces well the faint “fangs” seen in the northern part of the disk. If this structure is indeed part of the envelope, it suggests a young system at the end of the Class I phase. The serendipitous discovery of DraChi implies that such edge-on disks not associated with known star-forming regions (GoHam is another case) are still waiting to be discovered.
| Wavelength (m) | Flux (mJy) | Source | Instrument | Angular Resolution() |
| 0.23 | 0.14 0.1 | Bianchi et al. (2011) | GALEX | 5.3 |
| 0.48 | 6.46 0.20 | This Work | PS1 | 1.0 |
| 0.62 | 8.52 0.39 | This Work | PS1 | 1.0 |
| 0.75 | 9.71 0.37 | This Work | PS1 | 1.0 |
| 0.87 | 10.67 0.50 | This Work | PS1 | 1.0 |
| 0.96 | 12.41 0.80 | This Work | PS1 | 1.0 |
| 1.23 | 14.84 0.45 | Skrutskie et al. (2006) | 2MASS | 5 |
| 1.66 | 21.1 0.60 | Skrutskie et al. (2006) | 2MASS | 5 |
| 2.16 | 39.2 0.76 | Skrutskie et al. (2006) | 2MASS | 5 |
| 3.4 | 57.3 1.21 | Cutri et al. (2014) | WISE | 6.1 |
| 4.6 | 116 2.14 | Cutri et al. (2014) | WISE | 6.4 |
| 12 | 310 3.10 | Cutri et al. (2014) | WISE | 6.5 |
| 22 | 318 5.27 | Cutri et al. (2014) | WISE | 12.0 |
| 12 | 315 25 | Helou et al. (1988) | IRAS | 30 |
| 25 | 407 33 | Helou et al. (1988) | IRAS | 30 |
| 60 | 10700 86 | Helou et al. (1988) | IRAS | 60 |
| 100 | 14400 115 | Helou et al. (1988) | IRAS | 120 |
| 8.6 | 404 12 | Ishihara et al. (2010) | AKARI/IRC | 5.5 |
| 18.4 | 220 18 | Ishihara et al. (2010) | AKARI/IRC | 5.5 |
| 65 | 10840 570 | Kawada et al. (2007) | AKARI/FIS | 37 |
| 90 | 15490 276 | Kawada et al. (2007) | AKARI/FIS | 39 |
| 140 | 15440 1350 | Kawada et al. (2007) | AKARI/FIS | 58 |
| 160 | 12750 842 | Kawada et al. (2007) | AKARI/FIS | 61 |
Note. — Photometry presented in this paper from PS1 images and archival data available for this object.